OpenStudy (anonymous):

Find the solution of this equation : x^2+3x+2x+6 using the Fundamental Theorem of Algebra.

OpenStudy (anonymous):

this is actually not an equation you mean \(x^2+3x+2x+6=0\) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i guess you need to factor \[x^2+3x+2x+6=x(x+3)+2(x+3)=(x+2)(x+3)\] set each factor equal to zero and solve like \(x+2=0\iff x=-2\) for the first one

OpenStudy (anonymous):

just to be grumpy, you do not find anything using the fundamental theorem of algebra, so i am wondering who wrote this question

OpenStudy (anonymous):

A dumb study guide -.-

OpenStudy (anonymous):

yes, very dumb indeed